[figure: 9.7 cm]

MAGNETIC STRUCTURE OF THE FERRO/ANTIFERROMAGNET INTERFACE AND SPIN WAVE LOCALIZATION NEAR IT

A.G. Grechnev^{a} and A.S. Kovalev^{b} aKharkov State University,Svobody Sq.4, 310077, Kharkov, Ukraine, bInstitute for Low Temperature Physics and Engineering, 47 Lenin Ave., 310164, Kharkov, Ukraine

The multilayers of alternated ferro- and antiferromagnets are interesting and very promising objects in the area of technology. We consider theoretically a plane perfect interface between simple cubic ferromagnetic and antiferromagnetic halfspaces with an easy-axis anisotropy in the framework of a discrete classical Heisenberg model without taking into account magnetic dipole-dipole interactions. The classification of all possible collinear and canted magnetic structures of the interface is developed for both layered antiferromagnets and compensated interfaces. A phase diagram is presented for some possible collinear and noncollinear magnetic phases and the behavior of magnetization transformation under this collinear-canted transition is investigated. It is shown that the surface spin wave mode can exist at the ferro/antiferromagnetic interface and localized spin wave dispersion law has been calculated numerically.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : A.S.Kovalev

Address for correspondence : Institute for Low Temperature Physics and Engineering, 47 Lenin Ave., 310164 Kharkov Ukraine

Email address : kovalev@ilt.kharkov.ua